OpenStudy (anonymous):

If h(t) represents the height of an object above ground level at time t and h(t) is given by h(t)=−16t^2+14t+1. Find the height of the object at the time when the speed is zero.

OpenStudy (tyteen4a03):

Remember our ratio formula: speed = distance / time. (Another hint: since 0/anything = 0, basically you're solving for h(t) when t = 0.)

OpenStudy (anonymous):

the speed is the derivative of h(t) dh(t)/dt=-32t+14 -32t+14=0 hence t=14/32 plug t=14/32 in h(t) and you get 4,0625